# Constraining pseudo-Diracness with astrophysical neutrino flavors

### v2

@@ I Introduction
Here $\epsilon\sim 10^{-7}$ represents roughly the minimum value required to explain the observed baryon asymmetry. This provides further motivation to explore pseudo-Dirac mass splitting in the above range.
@@
The extragalactic origin of the HE neutrinos suggests that these particles travel, from their production to detection, a minimum distance of $L\approx 15$ kpc, equivalent to the radius of the Milky Way Galaxy. This estimate is notably conservative, considering that the Andromeda Galaxy, the closest major galaxy, lies 600 kpc away, while the nearest Active Galactic Nucleus (NGC5128) is situated roughly 4 Mpc from the Milky Way. This estimate is also compatible with the redshift evolution of BL Lac objects Ajello et al. (2014); Capanema et al. (2021). However, even with such a conservative choice of baseline, we anticipate that IceCube-Gen2 Aartsen et al. (2021a) should be able to impose the constraint $\delta m_{3}^{2}\lesssim 10^{-12}$ eV2 due to its improved flavor ratio measurements Abbasi et al. (2023b). This will leverage the constraint on $\delta m_{3}^{2}$ to the level of the current constraints on $\delta m_{1,2}^{2}$. On the other hand, in the more optimistic scenario where most potential neutrino sources reside farther away, for example at $z\sim 1$, as suggested by the star-formation rate Hopkins and Beacom (2006); Capanema et al. (2021), the bounds would improve: $\delta m_{3}^{2}\lesssim 10^{-17}$ eV2.
@@
In this work, we will use astrophysical neutrino flavor ratios to probe pseudo-Dirac neutrinos Crocker et al. (2000, 2002); Keranen et al. (2003); Beacom et al. (2004); Esmaili (2010); Shoemaker and Murase (2016); Brdar and Hansen (2019). In this approach, insensitivity to absolute neutrino flux is a double-edged sword. The advantage is that when there exist hierarchies between $\delta m_{i}^{2}$, stringent upper bounds on larger $\delta m_{i}^{2}$ can be derived as we will show in this work. The disadvantage is that if all $\delta m_{i}^{2}$ ($i=1,2,3$) are of the same order, we will lose the sensitivity. This is to be compared with studies which consider specific astrophysical sources where only certain ranges of $\delta m_{i}^{2}$ for the first few oscillations can be probed and, away from this region, the analysis is limited by absolute flux uncertainty De Gouvêa et al. (2020); Martinez-Soler et al. (2022); Rink and Sen (2024); Carloni et al. (2024); Dixit et al. (2024). If the flux from certain source can be determined quite accurately, and in particular, if all $\delta m_{i}^{2}$ are of the same order, this latter approach is superior. In the view that only very few astrophysical sources are identified, we will focus on flavor ratios approach and see how well this will fare. The paper is organized as follows: in Section II, we will review the parametrization of pseudo-Dirac model at low energy while in Section III, we will discuss the framework to calculate the oscillation probability in this scenario. Then, we will present our main results in Section IV and conclude in Section V.
@@ II Pseudo-Dirac model
${\cal M}$ can be diagonalized as ${\cal U}^{T}{\cal M}{\cal U}=\mathrm{diag}\left(\hat{m},\hat{m}\right)$ where the mass eigenstates are given by
@@
where $U_{0}$ is a $3\times 3$ unitary matrix while the rest of the $3\times 3$ matrices $A$, $B$, $C$ and $D$ are constrained by $\mathcal{U}\mathcal{U}^{\dagger}=\mathcal{U}^{\dagger}\mathcal{U}=I_{3\times 3}$. For example, from $\mathcal{U}\mathcal{U}^{\dagger}=I_{3\times 3}$, we have
@@
$AC^{\dagger}+BD^{\dagger}=CA^{\dagger}+DB^{\dagger}$ $=$ $0$ (20)
@@ III Oscillation probability
[…]
@@
By $\delta m_{k\neq j}^{2}=0$, we mean $\delta m_{k\neq j}^{2}\ll\delta m_{j}^{2}$ such that oscillations involving $\delta m_{k\neq j}^{2}$ have not developed.
@@ III.1 Neutrino flavor ratios
[…]
@@ IV Results
HE astrophysical neutrinos are generated in cosmic particle accelerators when accelerated protons collide with matter ($pp$ interactions) and radiation ($p\gamma$ interactions) Margolis et al. (1978); Stecker (1979); Mucke et al. (2000); Kelner et al. (2006); Hummer et al. (2010a). These collisions produce mesons, particularly charged pions ($\pi^{\pm}$), which then decay into neutrinos. The decay chain $\pi^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}\rightarrow e^{-}+\bar{\nu}_{e}+\nu_{\mu}+\bar{\nu}_{\mu}$, along with its charge conjugate reaction, produces two muon neutrinos for every electron neutrino: $(1,2,0)_{S}$, where the subscript $S$ denotes the initial proportions at the source [^3] Neutrinos and antineutrinos cannot be distinguished at the IceCube detector. The only exception is the Glashow Resonance Glashow (1960); Bhattacharya et al. (2011); Huang et al. (2023); Liu et al. (2023), but it suffers from a lack of statistics, with only one candidate event detected so far Aartsen et al. (2021b).. It is also possible that the decay chain described above is not complete, as the muon may lose energy due to synchrotron emission before decaying. As a result, this will generate the $\mu$-damped scenario, with the proportions $(0,1,0)_{S}$ Rachen and Meszaros (1998); Kashti and Waxman (2005); Kachelriess et al. (2008); Hummer et al. (2010b); Winter (2014). In this section, however, we assume neutrinos are produced only by the complete chain of $\pi^{\pm}$ decay and reach the Earth with averaged-out pseudo-Dirac oscillations, as described by eq. (37) and eq. (38), and compare these scenarios with the standard expectations for both $\pi$-decay and $\mu$-damped. More information regarding the $\mu$-damped production combined with the pseudo-Dirac scenario is reserved for the Appendix A.
@@
During propagation from the sources to the Earth, neutrino oscillations modify the initial flavor proportions. In the standard vacuum oscillation scenario, assuming the absence of matter effects Dev et al. (2023) and pseudo-Dirac mass splittings, the $\pi^{\pm}$ production channel results in flavor equipartition at Earth: $(1:2:0)_{S}\rightarrow(1:1:1)_{\oplus}$. To be more precise, in fig. 1 and fig. 2, we show the allowed region for the flavor composition at the Earth after considering the $3\sigma$ uncertainties of the oscillation parameters de Salas et al. (2021) where there is no perceptible distinction between normal and inverted mass orderings. Observe that the $\pi$-decay (blue region) best fit, in the standard case, is given by $(0.3,0.36,0.34)_{\oplus}$, which lies very close to the benchmark of total flavor equipartition. On the other hand, for the $\mu$-damped case (green region), the best fit is $(0.18,0.45.0.37)_{\oplus}$. Also represented in fig. 1 and fig. 2 are the regions corresponding neutrino flavor at the Earth in the presence of averaged-out pseudo-Dirac oscillation as described by eq. (37) and eq. (38) (red regions). We investigate all possible cases listed in the section III.
@@
Figure 2: Same as fig. 1, but for pseudo-Dirac scenarios named Case $jk$.
@@
We observe that Case 3 and Case 23 stand out as the most distinct from the standard scenarios, even when considering the current uncertainties in the oscillation parameters and production processes ($\pi$-decay or $\mu$-damped). What makes the regions corresponding to Cases 3 and 23 so distinct from the standard scenarios is that by decreasing the amount of $\nu_{3}$ in the flux, we automatically reduce the proportion of $f_{\mu,\oplus}$ (as well as $f_{\tau,\oplus}$, but in a smaller quantity), due to the larger mixing of this propagation eigenstate with $\nu_{\mu}$. Applying the same reasoning, if some of the $\nu_{1}$ eigenstates disappear (Case 1 or Case 12), $f_{e,\oplus}$ decreases, and the expected region shifts closer to the typical $\mu$-damped scenario. For Case 13, however, the decrease $f_{\mu\oplus}$ due to the disappearance of $\nu_{3}$ is balanced by the decrease in $f_{e\oplus}$ resulting from the disappearance of $\nu_{1}$, making the final region compatible with the standard $\pi$-decay. Moreover, $\nu_{2}$ has roughly equal mixing with all flavors, so its disappearance (Case 2) has a neutral effect. These results hold true for both normal and inverted mass orderings.
@@
Fig. 1 and fig. 2 also shows the $68\%$ and $95\%$ C.L. contours corresponding to the IceCube analysis Abbasi et al. (2022a) of the flavor composition of the HE astrophysical neutrinos using the High-Energy Starting Events (HESE) sample Abbasi et al. (2021). HESE is an all-flavor, all-sky search using 7.5 years of IceCube data for neutrinos with energies above 60 TeV. Observe that we cannot draw any conclusions from the current HESE bounds, which are still too weak to discriminate even between the standard scenarios. Nevertheless, future detectors, such as IceCube-Gen2 Aartsen et al. (2021a), with more statistics and improved capabilities, might significantly enhance the sensitivity to the flavor composition. In fig. 1 and fig. 2 we also show the $68\%$ and $99\%$ projected sensitivities of IceCube-Gen2 after 10 years of observations. The projection of IceCube-Gen2 is taken from Abbasi et al. (2023b) and assumes the best fit value to be $(0.3,0.36,0.34)_{\oplus}$, compatible with standard $\pi$-decay. Observe that with IceCube-Gen2, we can start to distinguish between the standard scenario and some pseudo-Dirac scenarios (Case 3 and Case 23), even if the production process ($\pi$-decay and/or $\mu$-damped) remains uncertain at that time. Furthermore, the more precise determination of several oscillation parameters in the next decade will contribute to astrophysical flavor measurements and enhance the discriminatory power of future detectors Song et al. (2021).

### v3

# Constraining the pseudo-Dirac nature of neutrinos using astrophysical neutrino flavor data

@@ I Introduction
Here $\epsilon\sim 10^{-7}$ represents roughly the minimum value required to explain the observed baryon asymmetry. Given current constraints, this can only be achieved by $\delta m_{3}^{2}$. This provides further motivation to explore the third pseudo-Dirac mass splitting within this range.
@@
The extragalactic origin of the HE neutrinos suggests that these particles travel, from their production to detection, a minimum distance of $L\approx 15$ kpc, equivalent to the radius of the Milky Way Galaxy. This estimate is notably conservative, considering that the Andromeda Galaxy, the closest major galaxy, lies 600 kpc away, while the nearest Active Galactic Nucleus (NGC5128) is situated roughly 4 Mpc from the Milky Way. This estimate is also compatible with the redshift evolution of BL Lac objects Ajello et al. (2014); Capanema et al. (2021). The important quantity which allows pseudo-Dirac neutrinos to be constrained is the pseudo-Dirac oscillating phase $\Phi_{j}\sim\delta m_{j}^{2}L/E$. If $\Phi_{j}\ll 1\implies\delta m_{j}^{2}\ll E/L$, the pseudo-Dirac pair remains coherent throughout the propagation and the oscillation approaches the standard scenario and we lose the sensitivity to constrain the scenario. If $\Phi_{j}\gg 1$, we have the decoherence among the pseudo-Dirac pair and since sterile states are not detected, the corresponding probability amplitude is suppressed by a factor of two. Even with a conservative choice of $L\sim 15$ kpc, under the assumption of hierarchical pseudo-Dirac mass splittings, we anticipate that IceCube-Gen2 Aartsen et al. (2021a) should be able to impose the constraint $\delta m_{3}^{2}\lesssim 10^{-12}$ eV2 due to its improved flavor ratio measurements Abbasi et al. (2023b). This will leverage the constraint on $\delta m_{3}^{2}$ to the level of the current constraints on $\delta m_{1,2}^{2}$. On the other hand, in the more optimistic scenario where most potential neutrino sources reside farther away, for example at $z\sim 1$, as suggested by the star-formation rate Hopkins and Beacom (2006); Capanema et al. (2021), the bounds would improve: $\delta m_{3}^{2}\lesssim 10^{-17}$ eV2.
@@
In this work, we will use astrophysical neutrino flavor ratios and the decoherence condition $\Phi_{j}\gg 1$ to probe pseudo-Dirac neutrinos Crocker et al. (2000, 2002); Keranen et al. (2003); Beacom et al. (2004); Esmaili (2010); Shoemaker and Murase (2016); Brdar and Hansen (2019). In this approach, insensitivity to absolute neutrino flux is a double-edged sword. The advantage is that when there exist hierarchies between $\delta m_{i}^{2}$, irrespective of spectral feature, capitalizing on the decoherence effect, stringent upper bounds on larger $\delta m_{i}^{2}$ can be derived as we will show in this work. The disadvantage is that if all $\delta m_{i}^{2}$ ($i=1,2,3$) are of the same order, we will lose the sensitivity since the transition probabilities for all flavors get approximately the same suppression. This is to be compared with studies which consider specific astrophysical sources where only certain ranges of $\delta m_{i}^{2}$ for the first few oscillations can be probed and, away from this region, the analysis is limited by absolute flux uncertainty and knowledge of the spectral feature De Gouvêa et al. (2020); Martinez-Soler et al. (2022); Rink and Sen (2024); Carloni et al. (2024); Dixit et al. (2024). If the flux from certain source can be determined quite accurately, and in particular, if all $\delta m_{i}^{2}$ are of the same order, this latter approach is superior. In the view that only very few astrophysical neutrino sources are identified, we will focus on flavor ratios approach and see how well this will fare.
The paper is organized as follows: in Section II, we will review the parametrization of pseudo-Dirac model at low energy while in Section III, we will discuss the framework to calculate the oscillation probability in this scenario. Then, we will present our main results in Section IV and conclude in Section V. In Appendix A, we investigate the effects of varying astrophysical neutrino sources.
@@ II Pseudo-Dirac model
[…]
@@
where $U_{0}$ is a $3\times 3$ unitary matrix while the rest of the $3\times 3$ matrices $A$, $B$, $C$ and $D$ are constrained by $\mathcal{U}\mathcal{U}^{\dagger}=\mathcal{U}^{\dagger}\mathcal{U}=I_{6\times 6}$. For example, from $\mathcal{U}\mathcal{U}^{\dagger}=I_{6\times 6}$, we have
@@
$AC^{\dagger}+BD^{\dagger}=CA^{\dagger}+DB^{\dagger}$ $=$ $0.$ (20)
@@ III Oscillation probability
$S$ $=$ $T\exp\left[-i\int_{0}^{t}dt^{\prime}H\left(t^{\prime}\right)\right],$ (25)
@@
By $\delta m_{k\neq j}^{2}=0$, we mean $\delta m^{2}_{k\neq j}\ll\delta m_{j}^{2}$ such that oscillations involving $\delta m_{k\neq j}^{2}$ have not developed. For example, considering the most distant identified source of astrophysical neutrinos PKS 1424+240 of 2.6 Gpc together with $E=1\,{\rm TeV}$ Paiano et al. (2017); Carloni et al. (2024), we need $\delta m^{2}_{k\neq j}\ll 10^{-20}\,{\rm eV}^{2}$ such that the oscillation involving $\delta m^{2}_{k\neq j}$ has not developed. Since very few astrophysical neutrino sources have been identified, the best we can currently do is to make assumptions based on the six possible scenarios mentioned above.
The relevant dataset of High-Energy Starting Events (HESE) Abbasi et al. (2021), from which neutrino flavors can be distinguished by IceCube, includes events with energies above 60 TeV and up to approximately 10 PeV. According to Eq. (35), the corresponding sensitivity range of pseudo-Dirac mass splittings for our analysis is shown for each case in Table 1. For example, in Case 3, where the expected regions for the standard and pseudo-Dirac scenarios do not overlap significantly (see next section), if the data remains consistent with the standard scenario at a given confidence level, we can place an upper bound on the mass splitting, $\delta m_{3}^{2}\lesssim 10^{-10}~\text{eV}^{2}$, assuming $\delta m_{1,2}^{2}=0$. This bound corresponds to the decoherence condition set by the upper end of the energy window, approximately 10 PeV. More stringent upper bounds, however, can be obtained by using lower-energy neutrinos around 100 TeV, for which the decoherence condition in Eq. (35) implies $\delta m_{3}^{2}\lesssim 10^{-12}~\text{eV}^{2}$. Statistically, this latter bound is expected to dominate (and is the one quoted in Table 1), as lower-energy neutrinos yield higher event statistics Abbasi et al. (2021).
Case Pseudo-Dirac mass splitting
Case 1 $\delta m_{1}^{2}\gtrsim 10^{-12}$ ${\rm eV}^{2}$, $\delta m_{2,3}^{2}=0$
Case 2 $\delta m_{2}^{2}\gtrsim 10^{-12}$ ${\rm eV}^{2}$, $\delta m_{1,3}^{2}=0$
Case 3 $\delta m_{3}^{2}\gtrsim 10^{-12}$ ${\rm eV}^{2}$, $\delta m_{1,2}^{2}=0$
Case 12 $\delta m_{1,2}^{2}\gtrsim 10^{-12}$ ${\rm eV}^{2}$, $\delta m_{3}^{2}=0$
Case 23 $\delta m_{2,3}^{2}\gtrsim 10^{-12}$ ${\rm eV}^{2}$, $\delta m_{1}^{2}=0$
Case 13 $\delta m_{1,3}^{2}\gtrsim 10^{-12}$ ${\rm eV}^{2}$, $\delta m_{2}^{2}=0$
Table 1: The cases with their corresponding pseudo-Dirac mass splitting ranges are valid for any astrophysical neutrinos with $E\sim 100\ {\rm TeV}$ from extragalactic sources i.e $L>15\,{\rm kpc}$. The meaning of $\delta m^{2}_{j}=0$ can be taken to be $\delta m^{2}_{j}\ll 10^{-20}\,{\rm eV}^{2}$ considering the most distant identified astrophysical neutrino source (see text for discussion).
@@ III.1 Neutrino flavor ratios
$f_{\beta,\oplus}^{j,\text{det}}=\frac{f_{\beta,\oplus}^{j}}{f_{e,\oplus}^{j}+f_{\mu,\oplus}^{j}+f_{\tau,\oplus}^{j}}\hskip 14.22636pt\text{and}\hskip 14.22636ptf_{\beta,\oplus}^{jk,\text{det}}=\frac{f_{\beta,\oplus}^{jk}}{f_{e,\oplus}^{jk}+f_{\mu,\oplus}^{jk}+f_{\tau,\oplus}^{jk}}.$ (41)
@@ IV Results
HE astrophysical neutrinos are generated in cosmic particle accelerators when accelerated protons collide with matter ($pp$ interactions) and radiation ($p\gamma$ interactions) Margolis et al. (1978); Stecker (1979); Mucke et al. (2000); Kelner et al. (2006); Hummer et al. (2010a). These collisions produce mesons, particularly charged pions ($\pi^{\pm}$), which then decay into neutrinos. The decay chain $\pi^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}\rightarrow e^{-}+\bar{\nu}_{e}+\nu_{\mu}+\bar{\nu}_{\mu}$, along with its charge conjugate reaction, produces two muon neutrinos for every electron neutrino: $(1,2,0)_{S}$, where the subscript $S$ denotes the initial proportions at the source [^3] Neutrinos and antineutrinos cannot be distinguished at the IceCube detector. The only exception is the Glashow Resonance Glashow (1960); Bhattacharya et al. (2011); Huang et al. (2023); Liu et al. (2023), but it suffers from a lack of statistics, with only one candidate event detected so far Aartsen et al. (2021b).. It is also possible that the decay chain described above is not complete, as the muon may lose energy due to synchrotron emission before decaying. As a result, this will generate the $\mu$-damped scenario, with the proportions $(0,1,0)_{S}$ Rachen and Meszaros (1998); Kashti and Waxman (2005); Kachelriess et al. (2008); Hummer et al. (2010b); Winter (2014). In this section, however, we assume neutrinos are produced only by the complete chain of $\pi^{\pm}$ decay and reach the Earth with averaged-out pseudo-Dirac oscillations, as described by eq. (37) and eq. (38), and compare these scenarios with the standard expectations for both $\pi$-decay and $\mu$-damped. More information regarding the $\mu$-damped production combined with the pseudo-Dirac scenario is reserved for the Supplemental Material.
@@
During propagation from the sources to the Earth, neutrino oscillations modify the initial flavor proportions. In the standard vacuum oscillation scenario, assuming the absence of matter effects Dev et al. (2023) and pseudo-Dirac mass splittings, the $\pi^{\pm}$ production channel results in flavor equipartition at Earth: $(1:2:0)_{S}\rightarrow(1:1:1)_{\oplus}$. To be more precise, in fig. 1 and fig. 2, we show the allowed region for the flavor composition at the Earth after considering the $3\sigma$ uncertainties of the oscillation parameters de Salas et al. (2021) where there is no perceptible distinction between normal and inverted mass orderings. Due to the smallness of $\sin\theta_{13}$, the effect of leptonic CP phase $\delta_{\rm CP}$ on allowed region is small and for this reason and to be as conservative as possible, we do not take into account the correlation between $\theta_{23}$ and $\delta_{\rm CP}$. Observe that the $\pi$-decay (blue region) using the best fit oscillation parameters of ref. de Salas et al. (2021), in the standard case, is given by $(0.3,0.36,0.34)_{\oplus}$, which lies very close to the benchmark of total flavor equipartition. On the other hand, for the $\mu$-damped case (green region), the best fit is $(0.18,0.45.0.37)_{\oplus}$. Also represented in fig. 1 and fig. 2 are the regions corresponding neutrino flavor at the Earth in the presence of averaged-out pseudo-Dirac oscillation as described by eq. (37) and eq. (38) (red regions). We investigate all possible cases listed in the section III.
@@
Figure 2: Same as fig. 1, but for pseudo-Dirac scenarios named Case jkjk.
@@
We observe that Case 3 and Case 23 stand out as the most distinct from the standard scenarios, even when considering the current uncertainties in the oscillation parameters and production processes ($\pi$-decay or $\mu$-damped). What makes the regions corresponding to Cases 3 and 23 so distinct from the standard scenarios is that by decreasing the amount of $\nu_{3}$ in the flux, we automatically reduce the proportion of $f_{\mu,\oplus}$ [^4] This fraction here should be undertood as the detected fraction $f^{\text{det}}$. (as well as $f_{\tau,\oplus}$, but in a smaller quantity), due to the larger mixing of this propagation eigenstate with $\nu_{\mu}$. Applying the same reasoning, if some of the $\nu_{1}$ eigenstates disappear (Case 1 or Case 12), $f_{e,\oplus}$ decreases, and the expected region shifts closer to the typical $\mu$-damped scenario. For Case 13, however, the decrease $f_{\mu,\oplus}$ due to the disappearance of $\nu_{3}$ is balanced by the decrease in $f_{e,\oplus}$ resulting from the disappearance of $\nu_{1}$, making the final region compatible with the standard $\pi$-decay. Moreover, $\nu_{2}$ has roughly equal mixing with all flavors, so its disappearance (Case 2) has a smaller effect. These results hold true for both normal and inverted mass orderings.
@@
Fig. 1 and fig. 2 also shows the $68\%$ and $95\%$ C.L. contours corresponding to the IceCube analysis Abbasi et al. (2022a) of the flavor composition of the HE astrophysical neutrinos using the HESE sample Abbasi et al. (2021). HESE is an all-flavor, all-sky search using 7.5 years of IceCube data for neutrinos with energies above 60 TeV. Observe that we cannot draw any conclusions from the current HESE bounds, which are still too weak to discriminate even between the standard scenarios. Nevertheless, future detectors, such as IceCube-Gen2 Aartsen et al. (2021a), with more statistics and improved capabilities, might significantly enhance the sensitivity to the flavor composition. In fig. 1 and fig. 2 we also show the $68\%$ and $99\%$ projected sensitivities of IceCube-Gen2 after 10 years of observations. The projection of IceCube-Gen2 is taken from Abbasi et al. (2023b) and assumes the best fit value to be $(0.3,0.36,0.34)_{\oplus}$, compatible with standard $\pi$-decay. Observe that with IceCube-Gen2, we can start to distinguish between the standard scenario and some pseudo-Dirac scenarios (Case 3 and Case 23), even if the production process ($\pi$-decay and/or $\mu$-damped) remains uncertain at that time. In Table 2, we summarize the best-fit flavor fractions for each of the cases illustrated in fig. 1 and fig. 2, normalized as in Eq. (41). In Appendix A, we show the results assuming $\mu$-damped channel and combination with $\pi$-decay and neutron decay channels of varying proportions for pseudo-Dirac scenario and in those cases, the differences between the standard and pseudo-Dirac scenarios are less pronounced.
At the Source
[…]
At the Earth
Scenario ($f_{e,\oplus}^{\text{det}}$, $f_{\mu,\oplus}^{\text{det}}$, $f_{\tau,\oplus}^{\text{det}}$)
Std. oscillations (0.30, 0.36, 0.34)
Case 1 (0.24, 0.40, 0.36)
Case 2 (0.29, 0.36, 0.35)
Case 3 (0.36, 0.31, 0.33)
Case 12 (0.22, 0.42, 0.36)
Case 13 (0.32, 0.35, 0.33)
Case 23 (0.38, 0.30, 0.32)
Table 2: Flavor composition at Earth for different scenarios using the best fit oscillation parameters of ref. de Salas et al. (2021), assuming pion decay at the source.
Finally, the more precise determination of several oscillation parameters in the next decade will contribute to astrophysical flavor measurements and enhance the discriminatory power of future detectors and taking into account the projected increase in precision in measurements of mixing parameters $\theta_{12}$ and $\theta_{23}$ in year 2040, the allowed regions practically shrink to best-fit points indicated by colored circles and triangle in fig. 1 and fig. 2 Song et al. (2021).